OpenStudy (anonymous):

The following curve passes through (3,1). Use the local linearization of the curve to find the approximate value of y at x=2.8.

OpenStudy (amistre64):

the following? curve?

OpenStudy (amistre64):

we need a curve in order to get the derivative, slope of the tangent line at 3,1 in order to approximate it with

OpenStudy (anonymous):

can we do it without a curve? like what if it just said a curve passes through (3,1).....

OpenStudy (amistre64):

no, becasue the linearization depends on the tangent line to the curve at that point; the point itself has no value to it alone

OpenStudy (anonymous):

oh the curve is 2x^2y+y=2x+13

OpenStudy (amistre64):

d/dx (2 x^2 y + y = 2x + 13) 4xy + 2x^2 y' + y' = 2 4xy + y'(2x^2 + 1) = 2 y' = (2- 4xy)/(2x^2 +1)

OpenStudy (amistre64):

that is the equation for the slope of the tangent line, now we use 3,1 to get the slopes value

OpenStudy (amistre64):

y' = (2- 4(3)(1))/(2(3)^2 +1) = 2 - 12 / 18 +1 = -10/19 = -5/8

OpenStudy (amistre64):

now we can use this in a linear equation for a slope Ty = -5x/8 +5(3)/8 +8/8 = -5x/8 +(15+8)/8 = -5x/8 +23/8 this will gives us an approximation at x=2.8 now

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

But, these are my posibble answer?? 0.5 0.9 0.95 1.1 1.4

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-5%282.8%29%2F8+%2B23%2F8 1.125 is close to 1.1

OpenStudy (anonymous):

got you